# WW-FL: Secure and Private Large-Scale Federated Learning

### v3

@@ 2. Preliminaries and Related Work
Hierarchical Federated Learning (HFL). Another approach for FL, called hierarchical FL (HFL) addresses scalability and heterogeneity issues in real-world systems (Bonawitz et al., 2019; Deng et al., 2021; Wang et al., 2021). In HFL, aggregation occurs at multiple levels, forming a hierarchical structure in which the aggregated values from one level serve as inputs for the above level. This procedure eventually reaches the top level, where the final model is aggregated (Deng et al., 2021; Wang et al., 2021). Previous works in the HFL setting have primarily focused on improving scalability and communication. However, (Yang, 2021) introduced methods for ensuring the privacy of individual client updates in HFL, allowing for secure aggregation. Despite this advancement, (Yang, 2021) did not address global model privacy or robustness against malicious users, which we provide in WW-FL.
@@
Poisoning & Robust Aggregation. One potential drawback of FL over PPML techniques is that FL increases the attack surface for malicious actors: in FL, each user trains their own model locally, allowing them to potentially manipulate their model (Biggio et al., 2012; Bhagoji et al., 2019). To counter these poisoning attacks, various robust aggregation methods have been proposed, in which local models are analyzed and appropriate measures such as outlier removal or norm scaling are applied (Yin et al., 2018; Li et al., 2019; Cao et al., 2021). See (Shejwalkar et al., 2022) for a comprehensive overview of different robust aggregations.
3. The WW-FL Framework
@@
[…]
$\mathsf{GS}_{i}\in\mathcal{G};i\in[\mathsf{N}_{\mathcal{G}}]$
Layer II: MPC Clusters
$\mathsf{CS}_{i}^{j}\in\mathcal{M}_{i};i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{M}_{i}}]$
Layer III: Clients
$\mathsf{C}_{i}^{j}\in\mathcal{C}_{i};i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{C}_{i}}]$
Figure 2. Three-layer architecture in WW-FL for federated training of a machine learning model.
Table 2. Notations used in our WW-FL framework.
@@
$\mathsf{N}_{\footnotesize\mathcal{M}}$ Total number of clusters.
@@
$\mathsf{CS}_{i}^{j}$ Layer II MPC Cluster Server; $\mathsf{CS}_{i}^{j}\in\mathcal{M}_{i}$. Here, $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{M}_{i}}]$
$\mathsf{C}_{i}^{j}$ Layer III Client; $\mathsf{C}_{i}^{j}\in\mathcal{C}_{i}$, $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{C}_{i}}]$
@@
3.1. WW-FL Architecture
Our WW-FL framework is based on a three-layer [^3] Additional layers can be easily incorporated depending on the scale of deployment. architecture and extends the established hierarchical FL (HFL) paradigm (Bonawitz et al., 2019; Wang et al., 2021; Yang, 2021). In HFL, clients are initially organized into clusters, and their local models are aggregated at the cluster level; these cluster-level models are then further aggregated globally, resulting in an additional layer of aggregation. We adopt the hierarchical approach in WW-FL for two main reasons: 1) to efficiently realize our distinct model training approach for global model privacy, and 2) to facilitate large-scale deployments across heterogeneous clients. However, there is one key difference: at the cluster level, WW-FL does not perform aggregation of models but private training on collective client data. This differs from both traditional and hierarchical FL since the data is not kept solely on the client devices. Nevertheless, WW-FL ensures that the data remains within the trusted domain of the client, allowing for accurate modeling of trust relationships between clients in real-world situations, such as trust among members of the same region or country (Marti and Garcia-Molina, 2006). Furthermore, this type of hierarchy is ubiquitous in real-world scenarios such as P2P gaming, organizations, and network infrastructures (Subramanian et al., 2002; Hong and Varghese, 2019; Lin et al., 2020).
@@ 3.1. WW-FL Architecture
Layer III: Clients. This layer is composed of $\mathsf{N}_{\footnotesize\mathcal{M}}$ distinct sets of clients $\mathcal{C}_{i}^{\sf U}$ (with $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]$), called clusters, which are formed based on specific criteria of the application (e.g., European Union (EU) member states for the EU smart metering scheme (Cuijpers and Koops, 2013; Commission, 2014)). Like standard FL, only a random subset of clients $\mathcal{C}_{i}\subseteq\mathcal{C}_{i}^{\sf U}$ will be selected by the training algorithm in an iteration $t\in[1,T]$.
@@
Layer II: MPC Clusters. The second layer consists of $\mathsf{N}_{\footnotesize\mathcal{M}}$ sets of distributed training servers $\mathcal{M}_{i}$ (with $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$), called MPC clusters, with each $\mathcal{M}_{i}$ corresponding to the cluster $\mathcal{C}_{i}^{\sf U}$ in Layer III. In iteration $t$, Layer I servers (denoted by $\mathcal{G}$) initiate ML training by sharing the current global model $W_{t-1}$ among the servers in $\mathcal{M}_{i}$. As will be discussed in §3.1, $W_{t-1}$ is also in a secret-shared form among $\mathcal{G}$, represented by ${\langle W_{t-1}\rangle}_{\mathcal{G}}$. To account for varying availability and trustworthiness of servers across regions, MPC clusters in WW-FL may use different MPC configurations and differ, e.g., in their corruption threshold and security model (Evans et al., 2018). Therefore, $\mathcal{G}$ uses the Reshare protocol (cf. Tab. 3) to convert the secret shares of ${\langle W_{t-1}\rangle}_{\mathcal{G}}$ to those of $\mathcal{M}_{i}$, i.e., ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}$.
@@
Given ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}$, the servers in $\mathcal{M}_{i}$ use the Train protocol (cf. Tab. 3) to run MPC-based PPML for private ML training (Knott et al., 2021; Keller and Sun, 2022) on the cumulative secret-shared data from all clients in the cluster $\mathcal{C}_{i}$, denoted by ${\langle{D}_{t}\rangle}_{\mathcal{M}_{i}}$. This data may include data from the same cluster in the previous iteration. Furthermore, by utilizing a larger pool of training data, we can leverage the known benefits of batching, resulting in faster convergence (Goyal et al., 2017; Bottou et al., 2018). After completing training, the servers in $\mathcal{M}_{i}$ run Reshare to secret-share the updated model with the Layer I servers, i.e., ${\langle W_{t}^{i}\rangle}_{\mathcal{G}}$.
@@
To preserve the system’s integrity, the servers for each MPC cluster must be chosen with care to ensure that clients are willing to share their data among the servers and that not all of the servers are colluding. One possible option is to build non-profit partnerships, such as in the MOC alliance (Zink et al., 2021), where organizations with mutual distrust can securely co-locate servers in the same data center with high-speed network connections. Alternatively, trusted entities like government organizations with limited infrastructure can host their servers in confidential cloud computing environments (Russinovich et al., 2021). If a client lacks trust in the WW-FL servers in its current cluster, the client can propose a new cluster with a partner that provides a non-colluding server.
@@
Layer I: Global Servers. The top layer consists of a set of MPC servers $\mathcal{G}$, named Global Servers, that securely aggregate trained models from all the MPC clusters in Layer II, similarly to a standard FL scheme with a distributed aggregator (Fereidooni et al., 2021; Ben-Itzhak et al., 2022; Rathee et al., 2023; Gehlhar et al., 2023). Given the locally trained models $W_{t}^{i}$ for $i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]$, the servers in $\mathcal{G}$ execute the secure aggregation protocol Agg  (Mansouri et al., 2023) to compute the updated global model in secret-shared form, i.e., ${\langle W_{t}\rangle}_{\mathcal{G}}$. The global servers in $\mathcal{G}$ use the Reshare protocol to distribute the aggregated model $W_{t}$ to each of the Layer II clusters $\mathcal{M}_{i}$ for the next iteration $t+1$.
@@
The distributed aggregator model, as outlined in §1, circumvents the possibility of several privacy attacks that were recently presented for single-server aggregators even when relying on maliciously secure aggregation protocols (So et al., 2021; Boenisch et al., 2021, 2022; Fowl et al., 2022; Wen et al., 2022; Zhao et al., 2023); for example, by excluding certain inputs from the computation and thus influencing the correctness (but not the security) of the protocol, a malicious single server can mount privacy attacks that allow to infer information about particular participants and their training data. In contrast, in our distributed aggregator setting, even when having only semi-honest protocol security, such attacks are detectable as long as at least one of the servers behaves honestly.
@@
0:  $\mathcal{G},\mathcal{M},\mathcal{C}$ $\mathcal{M}=\bigcup_{i}\mathcal{M}_{i}$, $\mathcal{C}=\bigcup_{i}\mathcal{C}_{i}$; $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$
@@
1:  initialize: ${\langle W_{0}\rangle}_{\mathcal{G}}\leftarrow\textsc{Share}(W_{0},\mathcal{G})$
@@
3:     for all $i\in\mathsf{N}_{\footnotesize\mathcal{M}}$ do # in parallel
4:        ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}\leftarrow\mathcal{G}.\textsc{Reshare}({\langle W_{t-1}\rangle}_{\mathcal{G}},\mathcal{M}_{i})$
5:        $\mathcal{C}_{i}\leftarrow\mathcal{M}_{i}.\textsc{Sample}(\mathcal{C}_{i}^{\sf U},t)$$\mathcal{C}_{i}^{\sf U}$ – total clients in $i$-th cluster
@@
7:           ${\langle{D}_{t}^{\mathsf{C}_{i}^{j}}\rangle}_{\mathcal{M}_{i}}\leftarrow{\mathsf{C}_{i}^{j}}.\textsc{Share}({D}_{t}^{\mathsf{C}_{i}^{j}},\mathcal{M}_{i})$
@@
[…]
10:        ${\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}}\leftarrow{\mathcal{M}_{i}}.\textsc{Train}({\langle W_{t-1}\rangle}_{\mathcal{M}_{i}},{\langle{D}_{t}\rangle}_{\mathcal{M}_{i}})$
11:        ${\langle W_{t}^{i}\rangle}_{\mathcal{G}}\leftarrow{\mathcal{M}_{i}}.\textsc{Reshare}({\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}},\mathcal{G})$
@@
13:     ${\langle W_{t}\rangle}_{\mathcal{G}}\leftarrow{\mathcal{G}}.\textsc{Agg}(\{{\langle W_{t}^{i}\rangle}_{\mathcal{G}}\}_{i\in[\mathsf{N}_{\footnotesize\mathcal{M}}]})$
@@
WW-FL – The Complete Picture. Alg. 1 (FL-Train) provides the training algorithm in WW-FL. We detail the necessary MPC functionalities along with the WW-FL training functionality $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ (Alg. 2) in §3.3. Although WW-FL has a 3-layer architecture, it can easily accommodate more levels of hierarchy depending on the size of the deployment. For this, additional layers of MPC clusters can be added between layers I and II, with the clusters performing secure aggregation instead of PPML training.
@@
Existing schemes for global model privacy, such as (Mandal and Gong, 2019) and (Fereidooni et al., 2021), only protect the model from either clients or aggregator servers, leaving the possibility of collusion especially in a cross-device setting. WW-FL addresses this issue by keeping the global model in secret-shared form, ensuring that no single entity or group of colluding entities (up to an allowed corruption threshold) can access the model. This provides a stronger notion of privacy and also protects against unauthorized use or misuse, such as a client disclosing the trained model to another organization for further training or commercial use. For scenarios demanding disclosure of the global model to authorized entities, such as consortium members, the global servers can utilize the Reveal protocol.
3.2. MPC Functionalities in WW-FL

### v4

@@ 2 Preliminaries and Related Work
Another approach for FL, called hierarchical FL (HFL) addresses scalability and heterogeneity issues in real-world systems [BEG+19, WWCJ22]. In HFL, aggregation occurs at multiple levels, forming a hierarchical structure in which the aggregated values from one level serve as inputs for the above level. This procedure eventually reaches the top level, where the final model is aggregated [DLR+21, WXL+21]. Most of the works in the HFL setting have primarily focused on improving scalability and communication [FHC+24, WHZ+25]. An exception is the work by [Yan21], which proposed methods to ensure the privacy of individual client updates in HFL, allowing for secure aggregation. However, their approach does not address global model privacy or provide robustness against malicious users, which we provide in WW-FL.
Poisoning & Robust Aggregation
@@
One potential drawback of FL over PPML techniques is that FL increases the attack surface for malicious actors: in FL, each user trains their own model locally, allowing them to manipulate their model [BCMC19, SH21]. To counter these poisoning attacks, various robust aggregation methods have been proposed, in which local models are analyzed and appropriate measures such as outlier removal or norm scaling are applied [YCRB18, LXC+19, CFLG21, BMP+24]. See [SHKR22, GGP24, ZBA+25] for a comprehensive overview of different robust aggregations.
Model Extraction Attacks
In model extraction attacks, an adversary leverages black-box access to a model’s outputs to deduce its structure without prior knowledge [LPLW24, ZBA+25]. While existing privacy-preserving FL frameworks using MPC or HE protect data privacy, they do not prevent such attacks, as the adversary still has access to model outputs. We consider model extraction prevention orthogonal to private inference in FL and suggest incorporating methods like rate limiting and differential privacy [ZLD+25] for future work.
3 The WW-FL Framework
@@
Table 2: Notations used in our WW-FL framework.
@@
$\mathsf{N}_{\scriptstyle\mathcal{M}}$ Total number of clusters.
@@
$\mathsf{CS}_{i}^{j}$ Layer II MPC Cluster Server; $\mathsf{CS}_{i}^{j}\in\mathcal{M}_{i}$. Here, $i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{M}_{i}}]$
$\mathsf{C}_{i}^{j}$ Layer III Client; $\mathsf{C}_{i}^{j}\in\mathcal{C}_{i}$, $i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{C}_{i}}]$
@@
[…]
$\mathsf{GS}_{i}\in\mathcal{G};i\in[\mathsf{N}_{\mathcal{G}}]$
Layer II: MPC Clusters
$\mathsf{CS}_{i}^{j}\in\mathcal{M}_{i};i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{M}_{i}}]$
Layer III: Clients
$\mathsf{C}_{i}^{j}\in\mathcal{C}_{i};i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}],j\in[\mathsf{N}_{\mathcal{C}_{i}}]$
Figure 2: Three-layer architecture in WW-FL for federated training of a machine learning model. The number of servers and clients shown is illustrative and not fixed. See §3.6 for further details.
3.1 WW-FL Architecture
Our WW-FL framework is based on a three-layer [^2] Additional MPC Cluster layers can be easily incorporated depending on the scale of deployment. architecture and extends the established hierarchical FL (HFL) paradigm [BEG+19, WXL+21, Yan21]. In HFL, clients are initially organized into clusters, and their local models are aggregated at the cluster level; these cluster-level models are then further aggregated globally, resulting in an additional layer of aggregation. We adopt the hierarchical approach in WW-FL for two main reasons: 1) to efficiently realize our distinct model training approach for global model privacy, and 2) to facilitate large-scale deployments across heterogeneous clients. However, there is one key difference: at the cluster level, WW-FL does not perform aggregation of models but private training on collective client data. This differs from both traditional and hierarchical FL since the data is not kept solely on the client devices. Nevertheless, WW-FL ensures that the data remains within the trusted domain of the client, allowing for accurate modeling of trust relationships between clients in real-world situations, such as trust among members of the same region or country [MG06]. Furthermore, this type of hierarchy is ubiquitous in real-world scenarios such as P2P gaming, organizations, and network infrastructures [SARK02, HV19, LSPJ20].
@@ 3.1 WW-FL Architecture
Layer III: Clients
This layer is composed of $\mathsf{N}_{\scriptstyle\mathcal{M}}$ distinct sets of clients $\mathcal{C}_{i}^{\sf U}$ (with $i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}]$), called clusters, which are formed based on specific criteria of the application (e.g., European Union (EU) member states for the EU smart metering scheme [CK13, Com14]). Like standard FL, only a random subset of clients $\mathcal{C}_{i}\subseteq\mathcal{C}_{i}^{\sf U}$ will be selected by the training algorithm in an iteration $t\in[1,T]$. This subset selection is performed by the MPC clusters in Layer II using the Sample algorithm, ensuring that the selection process adheres to the desired security guarantees.
Table 3: MPC functionalities used in WW-FL.
Algorithm Input(s) Description
Share D,V{D},{V} Generates secret shares of data D{D} as per target V{V}’s sharing semantics (${\langle{D}\rangle}_{{V}}$) with ${V}\in\{\mathcal{C},\mathcal{M}_{,}\mathcal{G}\}$.
Reshare ${\langle{D}\rangle}_{{U}},{V}$ Converts secret shares of data $\langle{D}\rangle$} from source U{U}’s sharing semantics to target V{V}.
Train ${\langle{W}\rangle},{\langle{D}\rangle}$ Performs PPML Training on ML model W{W} using the secret-shared data $\langle{D}\rangle$}.
Reveal ${\langle{D}\rangle}_{{U}},{V}$ Reconstructs secret-shared data $\langle{D}\rangle$} towards members in target set V{V}.
Sample $\mathcal{C}^{\sf U}$ Selects a subset of clients $\mathcal{C}$ from $\mathcal{C}^{\sf U}$.
@@
The second layer consists of $\mathsf{N}_{\scriptstyle\mathcal{M}}$ sets of distributed training servers $\mathcal{M}_{i}$ (with $i\in\mathsf{N}_{\scriptstyle\mathcal{M}}$), called MPC clusters, with each $\mathcal{M}_{i}$ corresponding to cluster $\mathcal{C}_{i}^{\sf U}$ in Layer III. In iteration $t$, Layer I servers (denoted by $\mathcal{G}$) initiate ML training by sharing the current global model $W_{t-1}$ among the servers in $\mathcal{M}_{i}$. As will be discussed in §3.1, $W_{t-1}$ is also in a secret-shared form among $\mathcal{G}$, represented by ${\langle W_{t-1}\rangle}_{\mathcal{G}}$. To account for varying availability and trustworthiness of servers across regions, MPC clusters in WW-FL may use different MPC configurations and differ, e.g., in their corruption threshold and security model [EKR18]. Therefore, $\mathcal{G}$ uses the Reshare protocol (cf. Tab. 3) to convert the secret shares of ${\langle W_{t-1}\rangle}_{\mathcal{G}}$ to those of $\mathcal{M}_{i}$, i.e., ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}$.
@@
Given ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}$, the servers in $\mathcal{M}_{i}$ use the Train protocol (cf. Tab. 3) to run MPC-based PPML for private ML training [KVH+21, KS22] on the cumulative secret-shared data from all clients in the cluster $\mathcal{C}_{i}$, denoted by ${\langle{D}_{t}\rangle}_{\mathcal{M}_{i}}$. This data may include data from the same cluster in the previous iteration. Furthermore, by utilizing a larger pool of training data, we can leverage the known benefits of batching, resulting in faster convergence [GDG+17, BCN18]. After completing training, the servers in $\mathcal{M}_{i}$ run Reshare to secret-share the updated model with the Layer I servers, i.e., ${\langle W_{t}^{i}\rangle}_{\mathcal{G}}$.
@@
To preserve the system’s integrity, the servers for each MPC cluster must be chosen with care to ensure clients are willing to share their data among the servers and that not all servers are colluding. One option is to build non-profit partnerships, such as in the MOC alliance [ZIC+21], where organizations with mutual distrust can securely co-locate servers in the same data center with high-speed network connections. Alternatively, trusted entities like government organizations with limited infrastructure can host their servers in confidential cloud computing environments [RCF+21]. If a client lacks trust in the WW-FL servers in its current cluster, the client can propose a new cluster with a partner that provides a non-colluding server.
Layer I: Global Servers
@@
The top layer consists of a set of MPC servers $\mathcal{G}$, named Global Servers, that securely aggregate trained models from all the MPC clusters in Layer II, similarly to a standard FL scheme with a distributed aggregator [FMM+21, RSWP23, GMS+23, BMP+24, TXLZ25]. Given the locally trained models $W_{t}^{i}$ for $i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}]$, the servers in $\mathcal{G}$ execute the secure aggregation protocol Agg  [MOJC23] to compute the updated global model in secret-shared form, i.e., ${\langle W_{t}\rangle}_{\mathcal{G}}$. The global servers in $\mathcal{G}$ use the Reshare protocol to distribute the aggregated model $W_{t}$ to each of the Layer II clusters $\mathcal{M}_{i}$ for the next iteration $t+1$.
@@
The distributed aggregator model, as outlined in §1, circumvents the possibility of several privacy attacks that were recently presented for single-server aggregators even when relying on maliciously secure aggregation protocols [BDS+22, FGC+22, WGF+22, SAG+23, BDS+23b, ZSE+24]; for example, by excluding certain inputs from the computation and thus influencing the correctness (but not the security) of the protocol, a malicious single server can mount privacy attacks that allow to infer information about particular participants and their training data. In contrast, in our distributed aggregator setting, even when having only semi-honest protocol security, such attacks are detectable as long as at least one of the servers behaves honestly.
@@
0:  $\mathcal{G},\mathcal{M},\mathcal{C}$ $\mathcal{M}=\bigcup_{i}\mathcal{M}_{i}$, $\mathcal{C}=\bigcup_{i}\mathcal{C}_{i}$; $i\in\mathsf{N}_{\scriptstyle\mathcal{M}}$
@@
1:  initialize: ${\langle W_{0}\rangle}_{\mathcal{G}}\leftarrow\textnormal{{Share}}(W_{0},\mathcal{G})$
@@
3:   for all $i\in\mathsf{N}_{\scriptstyle\mathcal{M}}$ do # in parallel
4:    ${\langle W_{t-1}\rangle}_{\mathcal{M}_{i}}\leftarrow\mathcal{G}.\textnormal{{Reshare}}({\langle W_{t-1}\rangle}_{\mathcal{G}},\mathcal{M}_{i})$# global servers $\mathcal{G}$ reshare $W_{t-1}$ with cluster servers
5:    $\mathcal{C}_{i}\leftarrow\mathcal{M}_{i}.\textnormal{{Sample}}(\mathcal{C}_{i}^{\sf U},t)$$\mathcal{C}_{i}^{\sf U}$ – total clients in $i$-th cluster
@@
7:     ${\langle{D}_{t}^{\mathsf{C}_{i}^{j}}\rangle}_{\mathcal{M}_{i}}\leftarrow{\mathsf{C}_{i}^{j}}.\textnormal{{Share}}({D}_{t}^{\mathsf{C}_{i}^{j}},\mathcal{M}_{i})$${D}_{t}^{\mathsf{C}_{i}^{j}}$ – client $\mathsf{C}_{i}^{j}$’s data in iteration $t$
@@
9:    ${\langle{D}_{t}\rangle}_{\mathcal{M}_{i}}\leftarrow\bigcup\limits_{j\in[\mathsf{N}_{\mathcal{C}_{i}}]}{\langle{D}_{t}^{\mathsf{C}_{i}^{j}}\rangle}_{\mathcal{M}_{i}}\bigcup{\langle{D}_{t-1}\rangle}_{\mathcal{M}_{i}}$${D}_{0}=\emptyset$
10:    ${\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}}\leftarrow{\mathcal{M}_{i}}.\textnormal{{Train}}({\langle W_{t-1}\rangle}_{\mathcal{M}_{i}},{\langle{D}_{t}\rangle}_{\mathcal{M}_{i}})$
11:    ${\langle W_{t}^{i}\rangle}_{\mathcal{G}}\leftarrow{\mathcal{M}_{i}}.\textnormal{{Reshare}}({\langle W_{t}^{i}\rangle}_{\mathcal{M}_{i}},\mathcal{G})$
@@
13:   ${\langle W_{t}\rangle}_{\mathcal{G}}\leftarrow{\mathcal{G}}.\textnormal{{Agg}}(\{{\langle W_{t}^{i}\rangle}_{\mathcal{G}}\}_{i\in[\mathsf{N}_{\scriptstyle\mathcal{M}}]})$
@@
Alg. 1 (FL-Train) provides the training algorithm in WW-FL. We detail the necessary MPC functionalities along with the WW-FL training functionality $\mathcal{F}_{\textsf{WW\mbox{-}FL}}$ (Alg. 3) in §3.3. Although WW-FL has a 3-layer architecture, it can easily accommodate more levels of hierarchy depending on the size of the deployment. For this, additional layers of MPC clusters can be added between layers I and II, with the clusters performing secure aggregation instead of PPML training.
@@
Existing schemes for global model privacy, such as [RSWP23] and [TXLZ25], only protect the model from either clients or aggregator servers, leaving the possibility of collusion especially in a cross-device setting. WW-FL addresses this issue by keeping the global model in secret-shared form, ensuring that no single entity or group of colluding entities (up to an allowed corruption threshold) can access the model. This provides a stronger notion of privacy and also protects against unauthorized use or misuse, such as a client disclosing the trained model to another organization for further training or commercial use. For scenarios demanding disclosure of the global model to authorized entities, such as consortium members, the global servers can utilize the Reveal protocol.
In §3.6, we provide a detailed explanation of the FL-Train algorithm for a scenario involving three Layer II MPC clusters, each using a different protocol: two-party [DSZ15], three-party [AFL+16, MR18], and four-party [KPRS22]. The global servers are instantiated using a two-party protocol with a helper, based on CrypTen [KVH+21].
3.2 MPC Functionalities in WW-FL